40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 17 01:55:16 2015, ericq, Update, LSC, Uncooperative AUX X 
    Reply  Fri Dec 18 17:37:04 2015, rana, Update, LSC, Uncooperative AUX X 
       Reply  Mon Dec 21 02:29:49 2015, ericq, Update, LSC, AUX X RIN measurements auxXRIN.pdf
          Reply  Tue Jan 5 02:54:38 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
             Reply  Tue Jan 5 13:17:06 2016, rana, Update, LSC, AUX X Freq Noise attempt 
                Reply  Tue Jan 5 16:33:45 2016, ericq, Update, LSC, AUX X Freq Noise attempt 
                   Reply  Thu Jan 7 04:28:39 2016, ericq, Update, LSC, AUX X Freq Noise measured PLLspec.pdfPLL_OLG.pdf
                      Reply  Thu Jan 7 16:52:32 2016, ericq, Update, LSC, AUX X Freq Noise measured AUXfreqnoise.pdfauxXmeasurements.zip
                         Reply  Thu Jan 7 19:04:25 2016, Koji, Update, LSC, AUX X Freq Noise measured 
                            Reply  Fri Jan 8 14:47:33 2016, ericq, Update, LSC, AUX Y Freq Noise measured AUX_freqnoises.pdfAUXY_Jan8.zip
                               Reply  Fri Jan 8 20:02:49 2016, rana, Update, LSC, AUX Y Freq Noise measured 
                               Reply  Sat Jan 9 00:39:15 2016, gautam, Update, LSC, AUX Y Freq Noise re-measured AuxPLL.pdf
                                  Reply  Mon Jan 11 19:01:56 2016, gautam, Update, LSC, PLL Marconi Investigation 6x
                                     Reply  Tue Jan 12 03:03:55 2016, ericq, Update, LSC, Frequently making noise 
                                        Reply  Tue Jan 12 19:38:31 2016, gautam, Update, LSC, Frequently making noise 2016_01_AUXLaser.tar.gzOLGs.pdfvariedCurrent.pdfvariedTemp.pdfPLL_setup.pdf
Message ID: 11924     Entry time: Sat Jan 9 00:39:15 2016     In reply to: 11921     Reply to this: 11925
Author: gautam 
Type: Update 
Category: LSC 
Subject: AUX Y Freq Noise re-measured 
Quote:

With the Y end laser, I was able to lock the PLL with a lower actuation range (1.6MHz/V), and with the PSL in both the free-running and MCL locked configurations.

I took spectra (attached) with the same actuation range (3.2 MHz/V) for the AUX X+PSL and AUX Y+PSL combinations (PSL shutter closed) just to keep things consistent. It looks like there is hardly any difference between the two combinations - could the apparent factor of 3 worse performance of the X end laser have been due to different actuation ranges on the Marconi? 

I've not managed to take a spectrum for the proposed replacement Lightwave laser on the PSL table, though with Eric's help, I've managed to find the beatnote (at a temperature of 53.0195 degrees). I had to do some minor alignment tweaking for this purpose on the PSL table - the only optics I touched were the ones in the pink beam path in attachments 1 and 2 in this elog (the setup used to make the measurement is also qualitatively similar to attachment 3 in the same elog, except for the fact that we are feeding back to the Marconi and not the laser - a detailed sketch with specific components used will be put up later). I'll try and measure the frequency noise of this laser as well over the weekend and put up some spectra. 

With regards to possibly switching out the Lightwave on the PSL table for the (faulty?) Innolight at the X end, I've verified the following:

  • The beam-height from the Lightwave on the mount it is currently sitting on is the same as that from the Innolight on the X end table.
  • There is sufficient space on the X end table to house the Lightwave laser+mount

It remains to characterize the beam coming out from the Lightwave laser and do a mode matching calculation to see if we can use the same optics currently in place (with slight rearrangement) to realize a satisfactory mode-matching solution - I've obtained a beam profiler to do this from Liyuan and have the software setup, but have yet to do the beam scan - the plan is to do this on the SP table, but we've put off moving the Lightwave laser off the PSL table until we (i) establish conclusively that the X end laser is malfunctioning and (ii) check the frequency nosie of the Lightwave relative to the Aux lasers currently at the ends. 

The area around the Marconi is in a little disarray at the moment with a bunch of cables, SR560s, analyzers etc - I didn't want to disconnect the measurement setup till we're done with it. I have however turned both IR beat PDs on the PSL table off, and have reconnected the Marconi output to the Frequency Generation Unit and have set the carrier back to 11.066209MHz, +13dBm. 

Attachment 1: AuxPLL.pdf  26 kB  | Hide | Hide all
AuxPLL.pdf
ELOG V3.1.3-